OpenStudy (anonymous):

I don't understand what you have to do in this question. So, can someone please help? The prices of different types of toy cars vary from $1.25 to $2.45 when tax is included. Dennis plans to spend 25% of the $20 he has saved to buy several toy cars. Which is the best estimate of the number of toy cars Dennis can buy with the money he plans to spend? A. 8 B. 5 C. 3 D. 1

OpenStudy (anonymous):

first you want to take 25% of $20: .25*20=5 then you want to divide this by the max a toy car can cost 5/2.45=2.04, so he can buy 2 of the most expensive toy cars now you want to divide 5 by the least expensive toy cars: 5/1.25=4 So he can buy 4 inexpensive toy cars or 2 expensive toy cars. Take the average of 2 and 4 to get the best estimate (2+4)/2=3 So the answer is 3